# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 5.1 Definitions
Case 3: $\text{dist}(v_{0},w)>\text{dist}(v_{0},v)$ Since $(v,w)$ is an arc, and we are in a shortest-path graph from $v_{0}$, the only possible distance relationship for $w$ to be greater than $v$’s distance is $\text{dist}(v_{0},w)=\text{dist}(v_{0},v)+1$. This implies $w\in R_{i+2}$. Since $u\in R_{i}$ and $w\in R_{i+2}$, there cannot be a direct arc $(u,w)$ in a shortest-path graph (as this would imply $\text{dist}(v_{0},w)=\text{dist}(v_{0},u)+1=i+1$, contradicting $\text{dist}(v_{0},w)=i+2)$. Thus, $(u,w)\notin G$. By definition, $w\in\text{ext}(u,v)$.
@@
In every exhaustive case, we were able to show that $w$ belongs to exactly one of the sets $\text{int}(u,v)$, $\text{ext}(u,v)$, or $\text{back}(v)$. Therefore, these three sets form a pairwise disjoint partition of $N^{+}(v)$.
Although they are not drawn, we can visualize back arcs in the example 5.1. Because these rooted neighborhoods are ordered around a minimum out-degree node, the back arcs would go from a later neighborhood to an earlier neighborhood.
@@
We have now laid the groundwork to discuss our approach. Suppose we have a node $u$ whose degree we want to control. We have a node $v$ that is an out-neighbor of $u$. We will divide $v$’s neighbors into two sets: those that are also neighbors of $u$ and those that are not neighbors of $u$. Now that we have defined the three ways we can divide degree-interior, exterior, and back arcs-we see how a node’s degree will be partitioned among those nodes. Each of those sets is disjoint because the rooted neighborhoods, the nodes are being sent to, are disjoint. Further, we will spend a lot of time excluding ourselves from the set of back arcs. Thus we will focus our energy on dividing a node $v$’s out-neighbors that come from a parent $u$ into interior out-neighbors within the same rooted neighborhood and exterior out-neighbors in the next rooted neighborhood.
@@
Because of the disjoint nature of our rooted neighborhoods, the interior and exterior neighbors are also disjoint. This is because while the interior neighbors deal with the first neighbor of a parent, the exterior neighbor deals with the second neighbor of that same parent. These neighborhoods are defined to be disjoint from one another, causing the interior and exterior neighborhoods to be disjoint. We can also say that for a given node $v$ and a parent $u$, in the absence of back arcs, that node’s entire out-degree can be represented by its interior and exterior arcs through that parent.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a rooted neighborhood’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in section 7 (Back Arcs).
@@
Next, we represent a data structure that is the combination of the things we have presented. This data structure should not be a surprise: It is just stating things that we know about distance being a total order, and what it does to rooted neighborhoods, and nodes within those rooted neighborhoods. That is, by having a total order, we have a universal ordering of both of these claims. Secondly, as the SSNC asks for another metric to compare nodes (like degree or lexicographic order), this data structure allows us to compare nodes within the rooted neighborhoods by their degrees as well. We also see that there are definitions of interior and exterior out-neighbors inherited from the graph itself.
@@ A Graph Level Order (GLOVER) on a directed graph $G=(V,E)$ can be defined as follows:
When we look at this data structure, it seems like a very simple set of things. The graph is partitioned by BFS distance from a minimum out-degree node, and further organizes vertices within each partition by degree. However, at the time of this literature review, there has been no work that has constructed such a partitioning for the Seymour Second Neighborhood Conjecture.
Novelty is not always in difficulty or complexity. Sometimes, it lies in seeing what the others overlooked and formalizing what opens a new path forward. The simplicity of the data structure emerges not from arbitrary design, but from the natural demands of the contradiction-based approach we took to the conjecture.
A central example of this is the distinction between interior and exterior neighbors. We are first able to split a node’s degree equation into these components. Then we prove that the interior degree always doubles. We will be left with a traversal algorithm through exterior nodes of the graph with a question of whether there is a node whose exterior degree doubles.
To illustrate the necessity of the GLOVER data structure, consider a directed graph where an arc $(z,w)$ exists such that the node $z$ belongs to a later rooted neighborhood than the node $w$. This makes the arc $(z,w)$ a back arc.
Example 5.2.
[…]
Figure 7: This example illustrates a scenario where a back arc $(z,w)$ exists from $R_{i+1}$ to $R_{i}$. Such arcs can lead to overlaps in the reachable sets and complicate the analysis based purely on the level-ordered structure.
Example 5.2 illustrates this example with three rooted neighborhoods and the back arc $(z,w)$ connecting the neighborhoods $R_{i+2}$ and $R_{i+1}$. A structured approach is necessary to show the relationship between these rooted neighborhoods and these back arcs and the Seymour Second Neighborhood Conjecture overall. In particular, a naive construction of the exterior paths would lead to situations where a node $w$ appears in an exterior set associated with a path ending at $z$, say $\text{ext}(y,z)$. This would make $w$ a neighbor of $z$, whereas $z$ appears to be later than $w$ in our conceptual ordering of the rooted neighborhood partitions.
The issue arises when an arc $(z,w)$ exists such that the node $z$ is further from the node $v_{0}$ than $w$ (i.e $\text{dist}(v_{0},z)<\text{dist}(v_{0}<w)$). This means the arc $(z,w)$ goes from a further node to a closer node. This is a clear distinction from interior neighbors, where all arcs are either interior or parallel. We can also see a third type of transitive triangle that arises in the GLOVER data structure from exterior neighbors in Example 5.3 through multiple forward arcs and a single parallel arc. Further, by distinguishing between these three types of transitive triangles, we can treat them differently. We see that the transitive triangles with only forward and parallel arcs are no problem for our data structure. It is only the ones with back arcs. In fact, the interior neighbors become building blocks for our data structure.
@@
It will seem like back arcs are being avoided in oriented graphs, like traditional transitive triangles were. What we will see in section 7 (Back Arcs) is that this is not the case. What we will see is that although they share the transitive triangle property with interior neighbors, they are more similar to exterior neighbors. These arcs are still neighbors to a node in a different rooted neighborhood. What makes back arcs simpler is that once we have shown the decreasing nature of these rooted neighborhoods, and interior degree doubling, the existence of a back arc will imply a degree doubling node. Thus, we are not avoiding back arcs because of their difficulty, but because we have to prove a lot of things before we can use it on the back arcs.
Example 5.3.
[…]
Figure 8: This example illustrates a third configuration of a transitive triangle: nodes $y$ and $w$ are located within region $R_{i+1}$, and node $z$ is in region $R_{i+2}$. The arcs are $(y,w)$, $(y,z)$, and $(w,z)$. Both $(y,z)$ and $(w,z)$ are forward-facing exterior arcs, meaning they originate from a ’closer’ region and terminate in a ’further’ region. This type of transitive triangle, where all arcs flow consistently with the distance metric, exemplifies a structure that is compatible with the GLOVER data structure, further distinguishing it from transitive triangles that include problematic ’back arcs’.
@@
This is the type of scenario, particularly the problematic nature of back arcs and their connection to transitive triangles, that highlighted the need for a more robust and organized framework. The GLOVER data structure directly addresses this by first partitioning the graph into a sequence of rooted neighborhoods about the minimum out-degree node $v_{0}$. Within this organized framework, we are able to differentiate between different types of transitive triangles, those involving interior arcs, and those involving back arcs. The critical properties that are enforced by the GLOVER data structure (e.g., by ensuring that rooted neighborhoods decrease in size, will be shown in Section 6), will allow us to transform these potentially problematic arcs into instances where the node $w$ has a second neighborhood larger than its first neighborhood. This will make $w$ a degree doubling node, consistent with the conjecture.
@@
This deep understanding, revealed through the GLOVER data structure, transforms what were previously intractable combinatorial hurdles—such as those posed by transitive triangles—into manageable components, thereby forging a clear path to a complete proof.
5.2 Proofs Using the GLOVER
@@
We must remember that we are still under the assumption that we are trying to prove the SSNC by contradiction. That is, we are supposing it is not true and investigating the properties that must be true in a graph. We are now ready to look again at the set-theoretic aspects of the SSNC. First, though, we need to understand what the set of interior and exterior elements represents.
@@
Interior Cover Lemma Let $G=(V,E)$ be an oriented graph and $u\in V(G)$. Suppose that for every $x\in N^{+}(u)$, there exists a $v\in N^{+}(u)$ such that $(u,v)\in E$ and $(v,x)\in E$. Then:
@@
Recall that $\text{int}(u,v)=N^{+}(u)\cap N^{+}(v)$ for any $v\in N^{+}(u)$, where $u$ us a parent of $v$.
First, let $x\in N^{+}(u)$. By the premise, there exists $v\in N^{+}(u)$ such that $(u,v)\in E$ and $(v,x)\in E$. Thus, $x\in N^{+}(v)$, and $x\in\text{int}(u,v)$. Hence, $N^{+}(u)\subseteq\bigcup_{v\in N^{+}(u)}\text{int}(u,v)$.
Conversely, if $x\in\bigcup_{v\in N^{+}(u)}\text{int}(u,v)$, then there exists $v\in N^{+}(u)$ such that $x\in\text{int}(u,v)$. Therefore, $x\in N^{+}(u)$, and $\bigcup_{v\in N^{+}(u)}\text{int}(u,v)\subseteq N^{+}(u)$.
@@
Lemma 5.2 (Interior Cover Lemma) demonstrates that if every child out-neighbor $x$ of a parent $u$ is the tail of a directed triangle with $u$ as the source, then the out-neighborhood of $u$ can be expressed as the union of the intersections of the out-neighborhoods of $u$ and it’s out-neighbors. This implies that the out-neighbors of $u$ share the workload of covering the out-neighborhood. Instead of $u$ needing to reach each of its out-neighbors directly, the existence of these transitive triangles ensures that each out-neighbor $x$ is also reached by another out-neighbor $v$. This effectively distributes the responsibility of reaching out-neighbors among the nodes in $N^{+}(u)$, preventing any single node from bearing the entire burden. This distribution can help prevent the node $u$ from becoming a degree-doubling node, provided there are sufficient nodes in $N^{+}(u)$ to handle the distributed responsibility.
@@
Exterior Cover Lemma Let $G=(V,E)$ be a graph, and let $u\in V$. Suppose that for every vertex $x$ at distance two from $u$, there exists a $v\in N^{+}(u)$ such that $(v,x)\in E(G)$. Then:
@@
Recall that for any $v\in N^{+}(u)$, the exterior set is defined as:
@@
So $x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)$. Thus,
@@
Conversely, suppose $x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v)$. Then there exists $v\in N^{+}(u)$ such that $x\in\text{ext}(u,v)$, which means:
@@
What’s happening is that we have reformed the conjecture without even introducing the Decreasing Neighborhood Sequence Property (DNSP). Rather than talking about second out-neighbors being greater than first, these lemmas 5.2 and 5.3 are saying that in order for the Seymour Second Neighborhood Conjecture to fail at a node, we need to have load balancing. This means, first, that we cannot have an out-neighbor with 0 out-degree in the first out-neighborhood. More than that, though, it means that we must have cycles present in that rooted neighborhood.
Next, we will prove that any oriented graph can be represented in a GLOVER. We will need to demonstrate how the graph can be systematically partitioned into rooted neighborhoods while preserving the GLOVER properties. This includes maintaining a total order of the rooted neighborhood, allowing for the comparison of any two nodes, establishing a total order of those nodes within the rooted neighborhood, and establishing interior and exterior out-neighbors.
@@
GLOVER Representation of Oriented Graphs without Back Arcs Given an oriented graph $G=(V,E)$ without back arcs, $G$ can be represented by a GLOVER data structure.
@@
We aim to show that $G$ can be systematically partitioned into rooted neighborhoods that satisfy the key properties required by the GLOVER structure:
We select a node $v_{0}\in V$ of minimum out-degree. This node will serve as the root for defining rooted neighborhoods based on graph distance.
Define rooted neighborhoods $R_{i}=\{u\in V|\text{dist}(v_{0},u)=i\}$ for $i=0,1,2,\ldots,k$. Here, $\text{dist}(v_{0},u)$ is the length of the shortest directed path from $v_{0}$ to $u$. Since $v_{0}$ is in $R_{0}$, each $R_{i}$ collects all nodes at distance exactly $i$ from $v_{0}$.
@@
For $u\in R$
Interior neighbors of $v\in N^{+}(u)\subseteq R_{i+1}$
@@
Specifically, if $(v_{1},v_{2})\in E$ with $v_{1},v_{2}\in R_{i+1}$, and both are children of $u$, then $v_{2}$ is an interior neighbor of $v_{1}$ w.r.t. $u$.
Exterior neighbors of $v\in N^{+}(u)$:
@@
Nodes reachable from $u$ in two steps but not directly from $u$, and which are reachable from $v$:
$\text{ext}(u,v)=N^{++}(u)\cap N^{+}(v)$ where $N^{++}(u)$ are the nodes at distance two from $u$. These sets fulfill the structural properties required by the GLOVER definition, capturing the overlap and extension of rooted neighborhoods through first and second neighbors.
@@
The partition ${R_{0},R_{1},\ldots,R_{k}}$ forms a leveled hierarchy of rooted neighborhoods, with a total order on rooted neighborhoods and on nodes within each neighborhood.
Interior and exterior neighborhoods are well-defined, preserving the GLOVER properties.
Hence, under the assumption of no back arcs, $G$ can be represented by a GLOVER data structure.
This completes the proof.
@@
What the GLOVER data structure does is transform the oriented graph into an ordering of rooted neighborhoods with metrics. This ordering is more than just representative. The rooted neighborhoods help to organize the nodes. What this suggests is that each partition, or rooted neighborhood, is totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in the 2-dimensional plane. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$. Now, each rooted neighborhood has a proper place, with its nodes inside it. These rooted neighborhoods are similar to hyper-nodes in that context since they were defined as induced subgraphs based on the shortest path. In this context, the exterior edges act as hyper-edges between the rooted neighborhoods.
@@
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each rooted neighborhood maintains its place in the total ordering. It is a bold assumption to claim that oriented graphs have no back arcs, though. Oriented graphs come in all shapes and sizes; surely some exist with back arcs. As we will cover in section 7 (Back Arcs), the disjoint nature of the rooted neighborhoods tells us that if any two neighborhoods are ever not disjoint, this is both a necessary and sufficient condition for the existence of a back arc.

### v12

@@ 5.1 Definitions
Case 3: $\text{dist}(v_{0},w)>\text{dist}(v_{0},v)$ Since $v\to w\in G$ is an arc, and we are in a shortest-path graph from $v_{0}$, the only possible distance relationship for $w$ to be greater than $v$’s distance is $\text{dist}(v_{0},w)=\text{dist}(v_{0},v)+1$. This implies $w\in R_{i+2}$. Since $u\in R_{i}$ and $w\in R_{i+2}$, there cannot be a direct arc $u\to w\in G$ in a shortest-path graph (as this would imply $\text{dist}(v_{0},w)=\text{dist}(v_{0},u)+1=i+1$, contradicting $\text{dist}(v_{0},w)=i+2)$. Thus, $u\to w\notin G$. By definition, $w\in\text{ext}(u,v)$.
@@
In every exhaustive case, we were able to show that $w$ belongs to exactly one of the sets $\text{int}(u,v)$, $\text{ext}(u,v)$, or $\text{back}(v)$. Therefore, these three sets form a pairwise disjoint partition of $N^{+}(v)$. ∎
Lemma 5.1 (Partition of Node’s Degree) is showing that these interior neighbors, exterior neighbors, and back arcs are fundamentally linked in a much stronger way that first and second neighbors are not. That is, they form a partition of a node’s degree. With that, any question about the redundancy of these definitions should be erased. Secondly, we need to understand that this lemma is dissecting the child node’s degree of an arc $u\to v$, where the node $u$ belongs to an earlier neighborhood. The node partition lemma shows that the node $v$ can now have its degree partitioned into interior neighbors, exterior neighbors, and back arcs. The node $u$ can have its degree partitioned as well, but it would need to be based on its own parents from previous rooted neighborhoods.
@@
We have now laid the groundwork to discuss our approach. Suppose we have a node $u$ whose degree we want to control. We have a node $v$ that is an out-neighbor of $u$. Now that we have defined the three ways we can divide degree-interior, exterior, and back arcs-we see how a node’s degree will be partitioned among those nodes. Each of those sets is disjoint because the rooted neighborhoods, the nodes are being sent to, are disjoint. Further, we will spend a lot of time excluding ourselves from the set of back arcs. We will divide $v$’s neighbors into the two remaining sets: those that are also neighbors of $u$ and those that are not neighbors of $u$. Thus we will focus our energy on dividing a node $v$’s out-neighbors that come from a parent $u$ into interior out-neighbors within the same rooted neighborhood and exterior out-neighbors in the next rooted neighborhood.
@@
Because of the disjoint nature of our rooted neighborhoods, the interior and exterior neighbors are also disjoint. This is because while the interior neighbors deal with the first neighbor of a parent, the exterior neighbor deals with the second neighbor of that same parent. These neighborhoods are defined to be disjoint from one another when there are no back arcs, causing the interior and exterior neighborhoods to be disjoint. We can also say that for a given node $v$ and a parent $u$, in the absence of back arcs, that node’s entire out-degree can be represented by its interior and exterior arcs through that parent.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a rooted neighborhood’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in section 7 (Back Arcs).
@@
Next, we represent a data structure that is the combination of the ideas we have presented. This data structure should not be a surprise: It is just stating facts that we know about distance being a total order, and what it does to rooted neighborhoods, and nodes within those rooted neighborhoods. That is, by having a total order, we have a universal ordering of both of these claims. Secondly, as the SSNC asks for another metric to compare nodes (like degree or lexicographic order), this data structure allows us to compare nodes within the rooted neighborhoods by their degrees as well. We also see that there are definitions of interior and exterior out-neighbors inherited from the graph itself.
@@ A Graph Level Order (GLOVER) on a directed graph $G=(V,E)$ can be defined as follows:
    Leveled Rooted Neighborhood Structure: The vertices of $V$ with minimum out-degree node $v_{0}$ are partitioned into levels of Rooted Neighborhoods $R_{1},R_{2},\ldots,R_{n}$, where $R_{i}=\{v\in G:\text{dist}(v_{0},v)=i\}$, and $\text{dist}(v_{0},v)$ is the shortest path from $v_{0}$ to $v$.
@@
The Graph Level Order (GLOVER) is a novel and essential data structure at the heart of our proof of the SSNC. Before this data structure even begins with a BFS-style partition, we have to ensure that, under the proof by contradiction assumption, these neighborhood sizes exhibit a Decreasing Neighborhood Sequence Property (DNSP). That is, each node must have a second neighborhood that is strictly smaller than its first neighborhood. It is only then that the data structure can proceed to partitioning the graph by Breadth-First Search (BFS) distance from a minimum out-degree node, and further organizing vertices within each partition by a second metric, such as degree. This is not simply an artifact of BFS. The absence of this critical DNSP step could directly lead to a degree-doubling node, and harm the construction of proof by contradiction.
Fundamentally, the Graph Level Order is far more than a simple sorting of nodes by BFS distance. Rather, in order to construct this data structure, we need both an understanding of the BFS algorithm, the DNSP, and how they impact the overall SSNC. The BFS algorithm simply determines a graph’s shortest path layers. The Graph Level Order gives those layers meaning, whether those meanings are trivially a degree doubling node, or the ability to dissect a node’s degree into core components: interior neighbors, exterior neighbors and back arcs, which was not possible with first and second neighbors. It gives nodes identifiers based on their distance from the root node. These identifiers allow us to dissect transitive triangles into six distinct types, so that instead of treating them uniformly and seeking graphs without transitive triangles, we can plan for each type. This data structure defines exterior neighbors so that we can now have trackers on Seymour diamonds, and they are no longer double counted second neighbors. It also defines rooted neighborhoods, and we can show that these decrease in size, which is the fundamental result we need to show that back arcs are not a problem for the conjecture.
This methodology, which extends foundational algorithmic concepts, is consistent with established practice in graph theory. Many previous authors have considered the BFS algorithm as a first step in research. Consider the Hopcroft-Karp algorithm for maximum matching [4]. This is very efficient and one of the most widely used algorithms for finding maximum matching today. The algorithm begins by conducting a BFS on the graph. Initially, algorithms would only add one path at a time, leading to a complexity of $O(m\cdot n)$. This algorithm came along and was built on top of BFS. Instead of doing one path at a time, it found all shortest paths in parallel and did them in one phase, cutting the run time to $O(\sqrt{n}\cdot m)$.
The Graph Level Order compares directly to this, as a central example of this is the distinction between interior, exterior neighbors, and back arcs. As we saw in Lemma 5.1 (Partition of Node’s Degree), we are first able to split a node’s degree equation into interior degree, exterior degree and back arc degree. We will then prove, using abstract algebra, that the interior degree of a node will double if it needs to map to $i$ other nodes. Next, we will prove that the size of the exterior neighbor sets will decrease along any path from the root node $v_{0}$ to any node on that path, as long as there are no back arcs. Then we will generalize the load balancing concept we have been talking about to show that every node $u_{i}\in R_{i}$ must have interior degree $i$.
This will leave us ready to develop a traversal algorithm through exterior nodes of the graph with a question of whether there is a node whose exterior degree doubles. This would not have been possible without this data structure which used the BFS layers to distinguish the distance from the minimum out-degree node as a critical missing link in the SSNC. We will also have to deal with back arcs, but as we will discuss in Section 7 (Back Arcs), that will not be a problem at that point because by then, we will be able to treat them more like exterior neighbors than interior neighbors. With the concepts of interior degree doubling and generalization of load balancing, we will be able to move on to back arcs.
@@
It will seem like back arcs are being avoided in oriented graphs, like traditional transitive triangles were by previous authors [2], [9]. What we will see in section 7 (Back Arcs) is that this is not the case. Although they share the transitive triangle property with interior neighbors, they are more similar to exterior neighbors. These arcs are still neighbors to a node in a different rooted neighborhood. What makes back arcs simpler is that once we have shown the decreasing nature of these rooted neighborhoods, and interior degree doubling, the existence of a back arc will imply a degree doubling node. Thus, we are not avoiding back arcs because of their difficulty, but because we have to prove a lot of concepts before we can use it on the back arcs.
@@
This is the type of scenario, particularly the problematic nature of back arcs and their connection to transitive triangles, that highlighted the need for a more robust and organized framework. The Graph Level Order directly addresses this by first partitioning the graph into a sequence of rooted neighborhoods about the minimum out-degree node $v_{0}$. Within this organized framework, we are able to differentiate between different types of transitive triangles, those involving interior arcs, and those involving back arcs. The critical properties that are enforced by the Graph Level Order (e.g., by ensuring that rooted neighborhoods decrease in size, will be shown in Section 6), will allow us to transform these potentially problematic arcs into instances where the node $w$ has a second neighborhood larger than its first neighborhood. This will make $w$ a degree doubling node, consistent with the conjecture.
@@
This deep understanding, revealed through the Graph Level Order, transforms what were previously intractable combinatorial hurdles—such as those posed by transitive triangles—into manageable components, thereby forging a clear path to a complete proof.
5.2 Proofs Using Graph Level Order
@@
We are trying to prove the SSNC by contradiction. That is, we are supposing it is not true and investigating the properties that must be true in a graph. We are now ready to look again at the set-theoretic aspects of the SSNC. First, though, we need to understand what the set of interior and exterior elements represents.
@@
Interior Cover Lemma Let $G$ be an oriented graph and $u\in G$. Suppose that for every $x\in N^{+}(u)$, there exists a $v\in N^{+}(u)$ such that $u\to v\in G$ and $v\to x\in G$. Then:
@@
Recall that
$\text{int}(u,v)=N^{+}(u)\cap N^{+}(v)$
for any $v\in N^{+}(u)$, where $u$ is a parent of $v$.
First, let $x\in N^{+}(u)$. By the premise, there exists $v\in N^{+}(u)$ such that $u\to v\in G$ and $v\to x\in G$. Thus, $x\in N^{+}(v)$, and $x\in\text{int}(u,v)$. Hence,
$N^{+}(u)\subseteq\bigcup_{v\in N^{+}(u)}\text{int}(u,v).$
Conversely, if
$x\in\bigcup_{v\in N^{+}(u)}\text{int}(u,v),$
then there exists $v\in N^{+}(u)$ such that
$x\in\text{int}(u,v).$
Therefore,
$x\in N^{+}(u),$
and
$\bigcup_{v\in N^{+}(u)}\text{int}(u,v)\subseteq N^{+}(u).$
@@
Lemma 5.2 (Interior Cover Lemma) demonstrates that if every child out-neighbor $x$ of a parent $u$ is the tail of a directed triangle with $u$ as the source, then the out-neighborhood of $u$ can be expressed as the union of the intersections of the out-neighborhoods of $u$ and its out-neighbors. This implies that the out-neighbors of $u$ share the workload of covering the out-neighborhood. Instead of $u$ needing to reach each of its out-neighbors directly, the existence of these transitive triangles ensures that each out-neighbor $x$ is also reached by another out-neighbor $v$. This effectively distributes the responsibility of reaching out-neighbors among the nodes in $N^{+}(u)$, preventing any single node from bearing the entire burden. This distribution can help prevent the node $u$ from becoming a degree-doubling node, provided there are sufficient nodes in $N^{+}(u)$ to handle the distributed responsibility.
@@
Exterior Cover Lemma Let $G$ be a graph, and let $u\in G$. Suppose that for every vertex $x$ at distance two from $u$, there exists a $v\in N^{+}(u)$ such that $v\to x\in G$. Then:
@@
Recall that for any $v\in N^{+}(u)$, the exterior set is defined as
@@
So
$x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v).$
Thus,
@@
Conversely, suppose
$x\in\bigcup_{v\in N^{+}(u)}\text{ext}(u,v).$
Then there exists $v\in N^{+}(u)$ such that $x\in\text{ext}(u,v)$, which means:
@@
What’s happening is that we have reformed the conjecture without even introducing the Decreasing Neighborhood Sequence Property (DNSP). Rather than talking about second out-neighbors being greater than first, these Lemmas 5.2 and 5.3 are saying that in order for the SSNC to fail at a node, we need to have load balancing. This means, first, that we cannot have an out-neighbor with 0 out-degree in the first out-neighborhood, i.e., we must have cycles present in that rooted neighborhood.
Next, we will prove that any oriented graph can be represented in a Graph Level Order. We will need to demonstrate how the graph can be systematically partitioned into rooted neighborhoods while preserving the Graph Level Order properties. This includes maintaining a total order of the rooted neighborhood, allowing for the comparison of any two nodes, establishing a total order of those nodes within the rooted neighborhood, and establishing interior and exterior out-neighbors.
@@
Graph Level Order Representation of Oriented Graphs without Back Arcs Given an oriented graph $G=(V,E)$ without back arcs, $G$ can be represented by a Graph Level Order.
@@
We aim to show that $G$ can be systematically partitioned into rooted neighborhoods that satisfy the key properties required by the Graph Level Order structure:
We select a node $v_{0}\in G$ of minimum out-degree. This node will serve as the root for defining rooted neighborhoods based on graph distance.
Define rooted neighborhoods $R_{i}=\{u\in G\mid\text{dist}(v_{0},u)=i\}$ for $i=0,1,2,\ldots,k$. Here, $\text{dist}(v_{0},u)$ is the length of the shortest directed path from $v_{0}$ to $u$. Since $v_{0}\in R_{0}$, each $R_{i}$ collects all nodes at distance exactly $i$ from $v_{0}$.
@@
For $u\in R_{i}$ Interior neighbors are defined for $v\in N^{+}(u)\subseteq R_{i+1}$
@@
Exterior neighbors are defined for $v\in N^{+}(u)$: Nodes reachable from $u$ in two steps but not directly from $u$, and which are reachable from $v$:
$\text{ext}(u,v)=N^{++}(u)\cap N^{+}(v)$
Consider an arc $u\to v\in G$, with $v\in R_{i+1}$. This makes $u$ a parent of $v$. Now consider the arc $v\to w\in G$, where $w\in N^{+}(v)$. There are three possible cases. Either $u\to w\in G$ or $u\to w\notin G,\text{ where }w\in R_{i}$ or $u\to w\notin G\text{ where }w\in R_{i+1}$.
Case 1: If $u\to w\in G$, then $w\in\text{int}(u,v)$.
Case 2: If $u\to w\notin G$, but $w\in R_{i+1}$, then $w\in\text{ext}(u,v)$.
Case 3: If $u\to w\notin G$ and $w\in R_{i+2}$, then $w\in\text{ext}(u,v)$.
@@
The partition ${R_{0},R_{1},\ldots,R_{k}}$ forms a leveled hierarchy of rooted neighborhoods, with a total order on rooted neighborhoods and on nodes within each neighborhood. Interior and exterior neighborhoods are well-defined, preserving the Graph Level Order properties. Hence, under the assumption of no back arcs, $G$ can be represented by a Graph Level Order.
This completes the proof. ∎
@@
What the Graph Level Order does is transform the oriented graph into an ordering of rooted neighborhoods with inner-metrics. This ordering is more than just representative. The rooted neighborhoods help to organize the nodes by giving them location identifiers. What this suggests is that each rooted neighborhood is totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in the 2-dimensional plane. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$. Now, each rooted neighborhood has a proper place, with its nodes inside it. These rooted neighborhoods are similar to hyper-nodes in that context since they were defined as induced subgraphs based on the shortest path. In this context, the exterior edges act as hyper-edges between the rooted neighborhoods.
With nodes now given these location identifiers, we can now think of the SSNC in terms of interior neighbors, exterior neighbors and back arcs instead of simply first and second neighbors. This is a more convenient thing because the new terms allow for a partition of a node’s degree (Lemma 5.1 (Partition of Node’s Degree)), whereas the older terms do not. Moreover, there have been problems with concepts like transitive triangles and double counting in the SSNC. We will show in the Lemma 5.2 (Classification of Transitive Triangles in the Graph Level Order) that by giving these nodes identifiers, we can distinguish different types of transitive triangles. This allows us to understand different ways to treat ’problematic’ transitive triangles (i.e. ones that have back arcs), vs ones that help us construct our data structure.
@@
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each rooted neighborhood maintains its place in the total ordering. It is a bold assumption to claim that oriented graphs have no back arcs, though. Oriented graphs come in all shapes and sizes. As we will cover in section 7 (Back Arcs), the disjoint nature of the rooted neighborhoods tells us that if any two neighborhoods are ever not disjoint, this is both a necessary and sufficient condition for the existence of a back arc.